A sample of 60 women is obtained, and their heights ( in inches ) and pulse rates ( in beats per minute ) are measured. The linear correlation coefficient is 0.234 and the equation of the regression line is ^y= 17.5 + 0.850x, where x represents height. The mean of the 60 heights is 63.4 in and the mean of the 60 pulse rates is 75.6 beats per minute. Find the best predicted pulse rate of a woman who is 74 inches tall. Use a significance level of a=0.01.
Answer #1

ŷ = 17.5 + 0.850x

Predicted value of y at x = 74

ŷ = 17.5 + (0.85) * 74 = 80.4

Best predicted pulse rate of a woman = 80.4 beats per minute